# A Database of Continued Fractions of Polynomial Type

### v2

@@ Abstract
We describe a database of 1400 continued fractions with polynomial coefficients, of which more than 1000 are new, both for interesting constants and for transcendental functions, and provide the database inside the  source of the paper. Look in particular at the section “Table of Contents”.
@@ 1 Introduction
    First, starting from the 200 or so CFs with polynomial coefficients (see below for the precise definition) existing in the literature, I have succeeded in creating more than 1000 additional ones, some of them I believe quite interesting.
@@
I have created a file containing these 1400 continued fractions together with their relevant information, which is readable from Pari/GP with the command V=readvec("file.gp"), if file.gp is the name given to the file (it is of course also readable by other programs). To obtain this file, from the arXiv repository from which you downloaded the pdf you must also obtain the  source, and the file is entirely contained between the last two comment commands. I hope that the arXiv moderators will not mind.
@@ 2.1 Computer Representation of Continued Fractions
written more compactly as $(a,b)=a(0)+b(0)/(a(1)+b(1)/(a(2)+\cdots))$. Note that this notation is different from that used by other authors such as [3]. I do not claim that my notation is better but simply that it is the one used in the present paper. We will always assume that $b(n)\neq 0$ (otherwise the CF terminates), and that $a(n)=0$ only for a finite number of $n$. We write as usual formally $p(n)/q(n)=a(0)+b(0)/(a(1)+b(1)/(a(2)+\cdots+b(n-1)/a(n)))$.
@@
The name polynomial is justified since it is easy to see that any CF with coefficients in $\mathbb{Q}(n)$ is equivalent (i.e., a CF with the same $p(n)/q(n)$) to one with coefficients in $\mathbb{Z}[n]$.
@@
$A$ has period 1 but $B$ has period 2, so we can represent the CF by
[[4,2*n+1],[[-4,9],[(2*n)^2,(2*n+3)^2]]]
@@ 2.2 The Database Formats
    Giving this to GP would make it choke (give an error message) because n^k is not a valid GP input when k is a variable. Thus we replace it in the file by
@@
    v[3] is itself a 4-component vector [d,[F,E,D,P],C,A] giving the exact speed of convergence of the CF, including a few terms of its asymptotic expansion.
@@
      v[3][2] gives the speed of convergence in our usual FEDP format, using d if necessary. The running example has v[3][2]=[0,1,0,1], meaning that we are in case $P^{+}$ with $P=1$. In some rare cases, the speed does not make any sense, in which case we set v[3][2]=0. In some other rare cases, the CF does not converge for real values of the variables (which is implicitly what is assumed), in which case we set v[3][2]=[0,0,0,0].
@@
      v[3][3] gives the constant $C$ such that $S-p(n)/q(n)\sim C/[FEDP]$, or 00 if unknown or has not been computed. Note that $C$ is always given as a closure, first because in some cases (as for the CF itself) it is not possible to do otherwise in GP, and second because writing C=()->Pi is both much clearer to read than C=3.1415... and also allows to have C to any accuracy.
@@
      v[3][4] gives the asympotic expansion A=1+..., such that $S-p(n)/q(n)=(C/[FEDP])A$, or 00 if unknown or not computed. The reserved letter x means $1/n$ and the reserved letter y means $1/n^{1/2}$.
@@
    The fact that the series begins at $n=0$ has been chosen for uniformity, but often leads to slightly more cumbersome formulas, for instance instead of $\zeta(2)=\sum_{n\geq 1}1/n^{2}$ we write instead $\zeta(2)=\sum_{n\geq 0}1/(n+1)^{2}$, hence represented by [0,1/(n+1)^2,1,1].
@@
    v[5] is used for testing. When empty, it means that the default values for the variables used in the CF can be used (presently $2/3$, $3/4$, $1/2$, $2$ for closures with arity less or equal to $4$). If it is a vector of values, these are taken instead of the default values. For instance for a closure of arity $2$, v[5]=[2/3,-1] means that one can take the default value $2/3$ for the first variable, but probably cannot use $3/4$ for the second, but $-1$ instead. Other values of v[5] have specific meanings which may change:
@@
    • v[5]=5: Choose $z=2i$, multiply by $-i$, test.
@@
    v[7] is a 2-component vector [A1,AD1] whose entries are strings describing the behavior of the CF under Apéry acceleration: A1 is for the Apéry diagonal process, and AD1 for the Apéry dual. The meanings of the strings are as follows: “NO” if Apéry is not applicable, “NOBA” if it is but the Bauer–Muir method introduces complicated denominators, “AP→label” (“APSI→label” if followed by a simplification) and “APD→label” if Apéry leads to the corresponding labeled CF, (“APD→SELF” it is self-dual), “APcomp” and “APDcomp” if the result is too complicated to be included in the encyclopedia so with no label, and a possible additional “sim” if similar to.
@@ 2.3 Table of Contents
The complete GP file of the 1400 CFs and definitions together with convergence and additional data can be obtained by obtaining the  source and saving the file contained between the comment statements which are at the end of the source file, and then reading it under GP with the command V=readvec("file.gp") if file.gp is the name that you gave to your saved file. The source file itself is divided into sections (included as comments so not read by GP) which helps to order the different CFs. The sections are the following, together with the range of corresponding indices of the vector V:
Definitions: 1–43
Constants: $2^{1/3}$: 44–77
Constants: $\pi$, $\log(2)$, and Periods of Degree 1: 78–182
Constants: $\pi^{2}$, $G$, and Periods of Degree 2: 183–273
Constants: $\pi^{3}$, $\zeta(3)$, and Periods of Degree 3: 274–301
Constants: $\pi^{4}$: 302–308
Constants: $\zeta(k)$ and $\pi^{k}$ for $k\geq 5$: 309–312
Constants: Linear Combinations of Zeta- and $L$-Values: 313–547
Miscellaneous Constants Related to $L$-Values: 548–552
Constants Related to Powers of $e=\exp(1)$: 553–590
Constants Involving Gamma Quotients: 591–665
Constants Coming from Bessel Functions: 666–681
Trivial Functions with an Arbitrary Sequence: 682–694
Other Trivial Functions: 695–705
Exponential Functions: 706–731
Logarithm Functions: 732–750
Power Functions: 751–775
Direct Hyperbolic and Trigonometric Functions: 776–841
Inverse Hyperbolic and Trigonometric Functions: 842–867
Gamma Function: One Variable: 868–909
Gamma Function: Several Variables: 910–923
Function $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$: 924–954
Function $\beta(z)=(\psi((1+z)/2)-\psi(z/2))/2$: 955–1030
Function $\psi^{\prime}(z)$: 1031–1054
Function $\beta_{1}(z)=-\beta^{\prime}(z)=\psi^{\prime}(z/2)/2-\psi^{\prime}(z)$: 1055–1065
Function $\psi^{\prime\prime}(z)$: 1066–1083
Function $\beta_{2}(z)=-\beta^{\prime}_{1}(z)=\psi^{\prime\prime}(z)-\psi^{\prime\prime}(z/2)/4$: 1084–1085
Function $\psi^{\prime\prime\prime}(z)$: 1086
Ordinary Generating Functions for Bernoulli Type Sequences: 1087–1211
Incomplete Gamma Functions: 1212–1223
Error Function and Related: 1224–1240
Jacobi Elliptic Functions: Laplace Transforms: 1241–1276
Complete Elliptic Integrals: 1277–1292
More Integrals: 1293–1315
Bessel and Related Functions: 1316–1338
General Hypergeometric Functions: ${}_{0}F_{1}$: 1339–1342
General Hypergeometric Functions: ${}_{1}F_{1}$: 1343–1350
General Hypergeometric Functions: ${}_{2}F_{0}$: 1351–1360
General Hypergeometric Functions: $U$: 1361–1362
General Hypergeometric Functions: ${}_{2}F_{1}$: 1363–1381
General Hypergeometric Functions: ${}_{2}F_{1}$ with $b=1$: 1382–1396
General Hypergeometric Functions: ${}_{3}F_{2}$ with $z=1$: 1397–1400
@@ 2.4 Beautifying the Continued Fractions
From a given CF in the database format, it is not difficult to create a “beautified” version which is more human-readable, and I advise the reader to do so. In fact, the author has done so and added a large quantity of additional information, thus creating a 500+ page book which may appear someday. In particular, some CFs may not be understandable without this additional information, but I believe that the present file is sufficiently useful as is.
Here is an example of beautification. Entries V[713] and V[714] are (newlines inserted to make the entries readable):
@@ 3 Some Remarks and Questions on the Database
    When the convergence is precisely $S-p(n)/q(n)\sim C/n^{k}$ with integral $k$, I do not know of a systematic way to compute the asymptotic expansion $A$ to more than $k-1$ terms, except when the CF comes from a sum of explicit rational functions. In many cases, this is due to the presence of more complicated terms in the asymptotic expansion, typically involving $\log(n)$. In all other convergence cases (except of course the logarithmic case), including $C/n^{k}$ when $k\notin\mathbb{Z}$, there is an algorithmic way to compute $A$.
@@
    The CFs 4.1.14.A0 for $\Gamma(1/3)^{3}/(2\pi/\sqrt{3})$ and 4.1.14.C for $\Gamma(2/3)^{3}/(2\pi/\sqrt{3})$ are extremely similar and have the same origin. They can both be Apéry accelerated giving complicated formulas, but the former has a very simple Apéry dual 4.1.14.B, while the latter has a very complicated Apéry dual which therefore we have not given. Is there an analogue of 4.1.14.B for $\Gamma(2/3)^{3}/(2\pi\sqrt{3})$

### v3

@@ Abstract
We describe a database of 1779 continued fractions with polynomial coefficients, of which more than 1500 are new, both for interesting constants and for transcendental functions, and provide the database inside the  source of the paper. Look in particular at the section “Table of Contents”.
@@ 1 Introduction
    First, starting from the 250 or so CFs with polynomial coefficients (see below for the precise definition) existing in the literature, I have succeeded in creating more than 1500 additional ones, some of them I believe quite interesting.
@@
I have created a file containing these 1779 continued fractions together with their relevant information, which is readable from Pari/GP with the command V=readvec("file.gp"), if file.gp is the name given to the file (it is of course also readable by other programs). To obtain this file, from the arXiv repository from which you downloaded the pdf you must also obtain the  source, and the file is entirely contained between the last two comment commands.
@@ 2.1 Computer Representation of Continued Fractions
written more compactly as $(a,b)=a(0)+b(0)/(a(1)+b(1)/(a(2)+\cdots))$. Note that this notation is different from that used by other authors such as [4]. I do not claim that my notation is better but simply that it is the one used in the present paper. We will always assume that $b(n)\neq 0$ for all $n$ (otherwise the CF terminates), and that $a(n)=0$ only for a finite number of $n$. We write as usual formally $p(n)/q(n)=a(0)+b(0)/(a(1)+b(1)/(a(2)+\cdots+b(n-1)/a(n)))$.
@@
The name polynomial is justified since it is clear that any CF with coefficients in $\mathbb{Q}(n)$ is equivalent (i.e., a CF with the same $p(n)/q(n)$) to one with coefficients in $\mathbb{Z}[n]$.
@@
AA has period 1 but $B$ has period 2, so we can represent the CF by
[[4,2*n+1],[[-4,9],[(2*n)ˆ2,(2*n+3)ˆ2]]]
@@ 2.2 The Database Formats
    Giving this to GP would make it choke (give an error message) because nˆk is not a valid GP input when k is a variable. Thus we replace it in the file by
@@
    v[3] is itself a 4-component vector [d,[F,E,D,P],C,A] giving the exact speed of convergence of the CF, including a few terms of its asymptotic expansion. If $L$ denotes the value of the CF and $p(n)/q(n)$ are the partial quotients, this notation means that
    $L-\dfrac{p(n)}{q(n)}=\dfrac{C}{n!^{F}E^{n}e^{\sqrt{Dn}}n^{P}}A\;,$
    with the following conventions:
@@
      v[3][2] gives the speed of convergence in the above FEDP format, using d if necessary. The running example has v[3][2]=[0,1,0,1], meaning that we are in case $P^{+}$ with $P=1$. In some rare cases, the speed does not make any sense, in which case we set v[3][2]=0. In some other rare cases, the CF does not converge for real values of the variables (which is implicitly what is assumed), in which case we set v[3][2]=[0,0,0,0].
@@
      v[3][3] gives the constant $C$ such that $S-p(n)/q(n)\sim C/[FEDP]$, or 0 if unknown or has not been computed. Note that $C$ is always given as a closure, first because in some cases (as for the CF itself) it is not possible to do otherwise in GP, and second because writing C=()->Pi is both much clearer to read than C=3.1415... and also allows to have C to any accuracy.
@@
      v[3][4] gives the asympotic expansion A=1+..., such that Sp(n)/q(n)=(C/[FEDP])AS-p(n)/q(n)=(C/[FEDP])A, or 0 if unknown or not computed. The reserved letter x means $1/n$ and the reserved letter y means $1/n^{1/2}$.
@@
    The fact that the series begins at $n=0$ has been chosen for uniformity, but often leads to slightly more cumbersome formulas, for instance instead of $\zeta(2)=\sum_{n\geq 1}1/n^{2}$ we write instead $\zeta(2)=\sum_{n\geq 0}1/(n+1)^{2}$, hence represented by [0,1/(n+1)ˆ2,1,1].
@@
    v[5] is used for testing. When empty, it means that the default values for the variables used in the CF can be used (presently $2/3$, $3/4$, $1/2$, 22 for closures with arity less or equal to $4$). If it is a vector of values, these are taken instead of the default values. For instance for a closure of arity $2$, v[5]=[2/3,-1] means that one can take the default value $2/3$ for the first variable, but probably cannot use $3/4$ for the second, but 1-1 instead. Other values of v[5] have specific meanings which may change:
@@
    • v[5]=5: Choose $z=2i$, multiply by i-i, test.
@@
    v[7] is a 2-component vector [A1,AD1] whose entries are strings describing the behavior of the CF under Apéry acceleration: A1 is for the Apéry diagonal process, and AD1 for the Apéry dual. The meanings of the strings are as follows: “NO” if Apéry is not applicable, “BAcomp” if it is but the Bauer–Muir method introduces complicated denominators, “AP→label” (“APSI→label” if followed by a simplification) and “APD→label” if Apéry leads to the corresponding labeled CF, (“APD→SELF” it is self-dual), “APcomp” and “APDcomp” if the result is too complicated to be included in the encyclopedia so with no label, and a possible additional “sim” if similar to.
@@ 2.3 Table of Contents
The complete GP file of the 1779 CFs and definitions together with convergence and additional data can be obtained by obtaining the  source and saving the file contained between the comment statements which are at the end of the source file, and then reading it under GP with the command V=readvec("file.gp") if file.gp is the name that you gave to your saved file. The source file itself is divided into sections (included as comments so not read by GP) which helps to order the different CFs. The sections are the following, together with the range of corresponding indices of the vector V:
Definitions: 1–45
Constants: $2^{1/3}$: 46–82
Constants: $\pi$, $\log(2)$, and Periods of Degree 1: 83–217
Constants: $\pi^{2}$, $G$, and Periods of Degree 2: 218–316
Constants: $\pi^{3}$, $\zeta(3)$, and Periods of Degree 3: 317–345
Constants: $\pi^{4}$: 346–352
Constants: $\zeta(k)$ and $\pi^{k}$ for $k\geq 5$: 353–356
Constants: Linear Combinations of Zeta- and $L$-Values: 357–633
Miscellaneous Constants Related to $L$-Values: 634–636
Constants Related to Powers of $e=\exp(1)$: 637–674
Constants Involving Gamma Quotients: 675–999
Constants Coming from Bessel Functions: 1000–1019
Trivial Functions with an Arbitrary Sequence: 1020–1030
Other Trivial Functions: 1031–1043
Exponential Functions: 1044–1071
Logarithm Functions: 1072–1096
Power Functions: 1097–1121
Direct Hyperbolic and Trigonometric Functions: 1122–1191
Inverse Hyperbolic and Trigonometric Functions: 1192–1220
Gamma Function: One Variable: 1221–1275
Gamma Function: Several Variables: 1276–1291
Function $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$: 1292–1324
Function $\beta(z)=(\psi((1+z)/2)-\psi(z/2))/2$: 1325–1400
Function $\psi^{\prime}(z)$: 1401–1426
Function $\beta_{1}(z)=-\beta^{\prime}(z)=\psi^{\prime}(z/2)/2-\psi^{\prime}(z)$: 1427–1437
Function $\psi^{\prime\prime}(z)$: 1438–1455
Function $\beta_{2}(z)=-\beta^{\prime}_{1}(z)=\psi^{\prime\prime}(z)-\psi^{\prime\prime}(z/2)/4$: 1456–1457
Function $\psi^{\prime\prime\prime}(z)$: 1458
Ordinary Generating Functions for Bernoulli Type Sequences: 1459–1583
Incomplete Gamma Functions: 1584–1595
Error Function and Related: 1596–1612
Jacobi Elliptic Functions: Laplace Transforms: 1613–1648
Complete Elliptic Integrals: 1649–1665
More Integrals: 1666–1688
Bessel and Related Functions: 1689–1714
General Hypergeometric Functions: $_{0}F_{1}$: 1715–1718
General Hypergeometric Functions: $_{1}F_{1}$: 1719–1726
General Hypergeometric Functions: $_{2}F_{0}$: 1727–1740
General Hypergeometric Functions: $U$: 1741–1742
General Hypergeometric Functions: $_{2}F_{1}$: 1743–1760
General Hypergeometric Functions: $_{2}F_{1}$ with $b=1$: 1761–1775
General Hypergeometric Functions: $_{3}F_{2}$ with $z=1$: 1776–1779
@@ 2.4 Beautifying the Continued Fractions
From a given CF in the database format, it is not difficult to create a “beautified” version which is more human-readable, and I advise the reader to do so. In fact, the author has done so and added a large quantity of additional information, thus creating a 630 page book which may appear someday. In particular, some CFs may not be understandable without this additional information, but I believe that the present file is sufficiently useful as is.
Here is an example of beautification. Entries V[1051] and V[1052] are (newlines inserted for better readability):
@@ 3 Some Remarks and Questions on the Database
    When the convergence is precisely $S-p(n)/q(n)\sim C/n^{k}$ with integral $k$, I do not know of a systematic way to compute the asymptotic expansion $A$ to more than k1k-1 terms, except when the CF comes from a sum of explicit rational functions. In many cases, this is due to the presence of more complicated terms in the asymptotic expansion, typically involving $\log(n)$. In all other convergence cases (except of course the logarithmic case), including $C/n^{k}$ when $k\notin\mathbb{Z}$, there is an algorithmic way to compute $A$.
@@
    The CFs 4.1.14.A0 for $\Gamma(1/3)^{2}/\Gamma(2/3)$ and 4.1.14.C for $\Gamma(2/3)^{2}/\Gamma(1/3)$ are extremely similar and have the same origin. They can both be Apéry accelerated giving complicated formulas, but the former has a very simple Apéry dual 4.1.14.B, while the latter has a very complicated Apéry dual which therefore we have not given. Is there an analogue of 4.1.14.B for $\Gamma(2/3)^{2}/\Gamma(1/3)$ ?
  10. 10.
    Similarly, there is a beautiful rapidly convergent CF 4.1.14.C6 as well as a very simple hypergeometric series for $\Gamma(2/3)^{2}/\Gamma(1/3)$, but apparently nothing of the sort for $\Gamma(1/3)^{2}/\Gamma(2/3)$.